I have tried everything but need desperate help!!!!
How do I graph these equations???
2x + y=-2
4x +2y=8
-2y = 3x + 1

Answer by JBarnum(2146)

first put them in y intercept form
y=mx+b
where m= slope and b= y intercept
y=-2x-2
m=-2/1
b=-2
on the graph plot (0,-2) from that point go down 2 and to the right 1 making point (1,-4) draw line
.
next
2y=-4x+8
y=-2x+4
m=-2/1
b=4
on the graph plot (0,4) from that point go down 2 and to the right 1 making point (1,2) draw line

there u go